The sum of the digits of a two-digit number is 7. The tens' digit is one more than twice the units' digit. Find the number.
Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)
Answer by greenestamps(13200)



If you just need an answer, look at pairs of digits whose sum is 7 and find the pair for which one of them is 2 more than two times the other.

If you are supposed to solve the problem as an exercise in algebra, then you have two basic options.

(1) Using two variables....

let t = tens digit
let u = units digit

Then

t+u = 7 (the sum of the two digits is 7)
t = 2u+1 (the tens digit is 1 more than twice the units digit)

Solve the pair of equations by any method you want. Substitution looks appropriate.

(2) Using a single variable....

let t = tens digit
then 7-t = units digit (because you are given t+u=7)

Then

t = 2(7-t)+1 (the tens digit is 1 more than twice the units digit)

Solve using basic algebra.